OpenStudy (anonymous):

a. Solve f=27.2 times m/ t

OpenStudy (anonymous):

can anyone help me? :/

OpenStudy (anonymous):

solve for what?

OpenStudy (ganpat):

it looks little incomplete to me..

OpenStudy (anonymous):

\[f=27.2\frac{m}{t}\] has \(f\) alone on one side, so at the moment it is "solved for"

OpenStudy (anonymous):

well • F is the force • m is the mass • t is the time in seconds im post to find this and its 27.2 x m divided by T

OpenStudy (ganpat):

well try this, little hypothetical.. Force = mass * acceleration (genrally g for freely following body).. therefore m * 9.81 = 27.2 * m /t t * 9.81 = 27.2 and t = 2.77 ... :-p.. i wildest guess...

OpenStudy (anonymous):

im so lost o.O i hate algebra. where in the world did u get 9.81? ughhhhhh

OpenStudy (ganpat):

ha ha, sorry its gravitational force.. g = 9. 81 m/s2.. m not sure on this answer, i mean u want this or not..